OpenStudy (anonymous):

What is the total amount that Matthew's bank will receive after lending him $8,000 for four years at an interest rate of 6 percent, compounded annually? total amount = P (1 + i)t

OpenStudy (anonymous):

P = 8000 i = 6/100 = 0.06 t = 4 \[Total = 8000(1.06)^4 = 10099.8\]